If (5,6) IS ON A LINE WITH THE SLOPE OF 4/3, GIVE THE COORDINATES FOR ANOTHER POINT ON THE LINE.
Answer by checkley71(8403)

Y=mX+b
6=4/3*5+b
6=20/3+b
b=6-20/3
b=(18-20)/3
b=-2/3 OR ANOTHER POINT IS THE Y INTERCEPT (0,-2/3)
#######################################################
MERRY CHRISTMAS